OpenStudy (anonymous):

Which of the systems below is equivalent to the given system? 0.01x - 0.05y = 0.1 0.4x + 0.3y = 0.9 10x- 5y= 10 and 4x+ 3y= 9 x- 5y= 10 and 4x+ 3y= 9 x- 5y= 1 and 4x+ 3y= 9

OpenStudy (lncognlto):

To solve this, you need to divide all values in the original equation by the coefficient of x, 0.01. What do we get then?

OpenStudy (anonymous):

no clue, I ain't too familiar with this, and my classes didn't get too specific in how to deal with them. sorry if I've inconvenienced you.

OpenStudy (lncognlto):

Not at all! We are here to help. Here's what to do:

OpenStudy (lncognlto):

0.01/0.01 = 1, -0.05/0.01 = 5, and 0.1/0.01 = 10. So x - 5y = 10

OpenStudy (anonymous):

multiplying the given eq both side by 100 u will get x- 5y= 10